If (29,1)is joined to (63,-16) in a co ordinate plane , then where will the line cut the x-axis ?
Answer by rfer(16322)

m=(-16-1)/(63-29)
m=-17/34
m=-1/2
----------------
y=mx+b
-16=-1/2*63+b
-16=-31.5+b
b=-16+31.5
b=15.5
----------------
y=-1/2x+15.5
-1/2x=-15.5
x=31
x intercept
(31,0)